A heavy object is dropped from a vertical height of 8.0 m. What is it’s speed when it hits the ground?
1
Expert's answer
2020-12-15T11:33:04-0500

We can find the speed of the object when it hits the ground from the law of conservation of energy:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot 9.8\ \dfrac{m}{s^2}\cdot 8\ m}=12.52\ \dfrac{m}{s^2}."

Answer:

"v=12.52\ \dfrac{m}{s^2}."